# UniMODE: Unified Monocular 3D Object Detection

### v1

@@ Abstract
Realizing unified monocular 3D object detection, including both indoor and outdoor scenes, holds great importance in applications like robot navigation. However, involving various scenarios of data to train models poses challenges due to their significantly different characteristics, e.g., diverse geometry properties and heterogeneous domain distributions. To address these challenges, we build a detector based on the bird’s-eye-view (BEV) detection paradigm, where the explicit feature projection is beneficial to addressing the geometry learning ambiguity when employing multiple scenarios of data to train detectors. Then, we split the classical BEV detection architecture into two stages and propose an uneven BEV grid design to handle the convergence instability caused by the aforementioned challenges. Moreover, we develop a sparse BEV feature projection strategy to reduce computational cost and a unified domain alignment method to handle heterogeneous domains. Combining these techniques, a unified detector UniMODE is derived, which surpasses the previous state-of-the-art on the challenging Omni3D dataset (a large-scale dataset including both indoor and outdoor scenes) by 4.9% $\rm AP_{3D}$, revealing the first successful generalization of a BEV detector to unified 3D object detection [^1] This paper has been accepted for publication in CVPR2024..
@@ 1 Introduction
The most critical challenge in unified 3D object detection lies in addressing the distinct characteristics of different scenarios. For example, indoor objects are smaller and closer in proximity, while outdoor detection needs to cover a vast perception range. Recently, Cube RCNN [5] has served as a predecessor in studying this problem. It directly produces 3D box predictions in the camera view and adopts a depth decoupling strategy to tackle the domain gap among scenes. However, we observe it suffers serious convergence difficulty and is prone to collapsing during training.
@@
First of all, as shown in Fig. 1 (a) and (b), the geometry properties (e.g., perception ranges, target positions) between indoor and outdoor scenes are quite diverse. Specifically, indoor objects are typically a few meters away from the camera, while outdoor targets can be more than 100m away. Since a unified BEV detector is required to recognize objects in all scenarios, the BEV feature has to cover the maximum possible perception range. Meanwhile, as indoor objects are often small, the BEV grid resolution for indoor detection needs to be precise. All these characteristics can lead to unstable convergence and significant computational burden. To address these challenges, we develop a two-stage detection architecture. In this architecture, the first stage produces initial target position estimation, and the second stage locates targets using this estimation as prior information, which helps stabilize the convergence process. Moreover, we introduce an innovative uneven BEV grid split strategy that expands the BEV space range while maintaining a manageable BEV grid size. Furthermore, a sparse BEV feature projection strategy is developed to reduce the projection computational cost by 82.6%.
@@
Combining all these innovative techniques, a Unified Monocular Object DEtector named UniMODE is developed, and it achieves state-of-the-art (SOTA) performance on the Omni3D benchmark. In the unified detection setting, UniMODE surpasses the SOTA detector, Cube RCNN, by an impressive 4.9% in terms of $\rm AP_{3D}$ (average precision based on 3D intersection over union). Furthermore, when evaluated in indoor and outdoor detection settings individually, UniMODE outperforms Cube RCNN by 11.9% and 9.1%, respectively. This work represents a pioneering effort to explore the generalization of BEV detection architectures to unified detection, seamlessly integrating indoor and outdoor scenes. It showcases the immense potential of BEV detection across a broad spectrum of scenarios and underscores the versatility of this technology.
@@ 2 Related Work
BEV detectors, on the other hand, transform image features from the 2D camera plane to the 3D physical space before generating results in 3D [12]. This approach benefits downstream tasks, as planning is also performed in the 3D space [18]. However, the challenge with BEV detectors is that the feature transformation process relies on accurate depth estimation, which can be difficult to achieve with only camera images [23]. As a result, convergence becomes unstable when dealing with diverse data scenarios [5].
@@ 3.1 Overall Framework
The feature head and depth head are responsible for projecting the image feature into the BEV plane and obtaining BEV feature. During this projection, we develop a technique to remove unnecessary projection points, which reduces the computing burden by about 82.6% (refer to Section 3.4). Besides, we propose the uneven BEV feature (refer to Section 3.3), which means the BEV grids closer to the camera enjoy more precise resolution, and the grids farther to the camera cover broader perception areas. This design well balances the grid size contradiction between indoor detection and outdoor detection without extra memory burden.
@@ 3.2 Two-Stage Detection Architecture
The geometry property difference is identified as an essential reason causing the unstable convergence of BEV detectors [15]. For example, the target position distribution difference makes it challenging for Transformer-based detectors to learn how to update the query reference points gradually toward concerned objects. In fact, through visualization, we find the reference point updating in the 6 Transformer decoders is disordered. As a result, if we adopt the classical deformable DETR architecture [41] to build a 3D object detector, the training is easy to collapse due to the inaccurate positions of learned reference points, resulting in sudden gradient vanishing or exploding.
@@
To overcome this challenge, we construct UniMODE in a two-stage detection fashion. In the first stage, we design a CenterNet [39] style head (the proposal head in Fig. 2) to produce detection proposals. Specifically, its predicted attributes include the 2D center Gaussion heatmap, offset from 2D centers to 3D centers, and 3D center depths of targets. The 3D center coordinates of proposals can be derived from these predicted attributes. Then, the proposals with top $M$ confidences are selected and encoded as $M$ proposal queries by an MLP layer. To account for any potential missed targets, another $N$ randomly initialized queries are concatenated with these proposal queries to perform information interaction in the 6 decoders of the second stage (the Transformer stage). In this way, the initial query reference points of the second detection stage are adjusted adaptively. Our experiments reveal that this two-stage architecture is essential for stable convergence.
@@ 3.3 Uneven BEV Grid
A notable difference between indoor and outdoor 3D object detection lies in the geometry information (e.g., scale, proximity) of objects to the camera during data collection. Indoor environments typically feature smaller objects located closer to the camera, whereas outdoor environments involve larger objects positioned at greater distances. Furthermore, outdoor 3D object detectors must account for a wider perceptual range of the environment. Consequently, existing indoor 3D object detectors typically use smaller voxel or pillar sizes. For instance, the voxel size of CAGroup3D [31], a SOTA indoor 3D object detector, is 0.04 meters, and the maximum target depth in the SUN-RGBD dataset [29], a classic indoor dataset, is approximately 8 meters. In contrast, outdoor datasets exhibit much larger perception ranges. For example, the commonly used outdoor detection dataset KITTI [8] has a maximum depth range of 100 meters. Due to this vast perception range and limited computing resources, outdoor detectors employ larger BEV grid sizes, e.g. the BEV grid size in BEVDepth [11], a state-of-the-art outdoor 3D object detector, is 0.8 meters.
@@ 3.5 Unified Domain Alignment
Class alignment loss. In the loss view, we aim to address the heterogeneous label conflict when combining multiple data sources. Specifically, there are 6 independently labeled sub-datasets in Omni3D, and their label spaces are different. For example, as presented in Fig. 4, although the Window class is annotated in ARKitScenes, while it is not labeled in Hypersim. As the label space of Omni3D is the union of all classes in all subsets, the unlabeled window in Fig. 4 (a) becomes a missing target that harms convergence stability.
@@
The two-stage detection architecture described in Section 3.2 can alleviate the aforementioned problem to some extent, because it helps the detector concentrate on foreground objects, and the unlabeled objects are overlooked to compute loss. To address this problem further, we devise a simple strategy, the class alignment loss. Specifically, denoting the label space of the $i_{\rm th}$ dataset as $\Omega_{i}$, we compute loss on the $i_{\rm th}$ dataset as:
@@ 4 Experiment
Implementation details. The perception ranges in the X-axis, Y-axis, and Z-axis of the camera coordinate system are $(-30,30)$, $(-40,40)$, $(0,80)$ meters, respectively. If without a special statement, the BEV grid resolution is $(60,80)$. The factor $\gamma$ defined in the class alignment loss is set to 0.2. $M$ and $N$ are set to 100. The adopted optimizer is AdamW, and the learning rate is set to $12e^{-4}$ for a batch size of 192. The experiments are primarily conducted on 4 A100 GPUs. The total loss includes two parts, the proposal head loss and the query FFN loss. The proposal head loss consists of the heatmap classification loss and depth regression loss. The query FFN loss comprises the classification loss (a cross-entropy loss) and regression loss ($L_{1}$ loss for predicting 3D center, dimension, and orientation). The total loss is the weighted sum of these loss items. No special loss is set for the $M$ proposal query generation.
Dataset. The experiments in this section are performed on Omni3D, the sole large-scale 3D object detection benchmark encompassing both indoor and outdoor scenes. Omni3D is built upon six well-known datasets including KITTI [8], SUN-RGBD [29], ARKitScenes [3], Objectron [1], nuScenes [6], and Hypersim [27]. Among these datasets, KITTI and nuScenes focus on urban driving scenes, which are real-world outdoor scenarios. SUN-RGBD, ARKitScenes, and Objectron primarily pertain to real-world indoor environments. Compared with outdoor datasets, the required perception ranges of indoor datasets are smaller and the object categories are more diverse. Hypersim, distinct from the aforementioned five datasets, is a virtually synthesized dataset. Thus, Hypersim allows for the annotation of object classes that are challenging to label in real scenes, such as transparent objects (e.g., windows) and very thin objects (e.g., carpets). The Omni3D dataset comprises a total of 98 object categories and 3 million 3D box annotations, spanning 234,000 images. The evaluation metric is $\rm AP_{3D}$, which reflects the 3D Intersection over Union (IoU) between 3D box prediction and label.
@@ 4.1 Performance Comparison
According to the results, we can observe that UniMODE achieves the best results in all metrics. It surpasses the SOTA Cube RCNN by 4.9% given the primary metric $\rm AP_{3D}$. Besides DLA34, we also try another backbone ConvNext-Base. This is because previous papers suggest that DLA34 is commonly used in camera-view detectors like Cube RCNN but not suitable for BEV detectors [13]. Since UniMODE is a BEV detector, only testing the performance of UniMODE with DLA34 is unfair. Thus, we also test the UniMODE with ConvNext-Base, and the result suggests that the performance is boosted significantly. Additionally, the speed of UniMODE is also promising. Test on 1 A100 GPU, the inference speeds of UniMODE under the high and low computing resource settings are 21.41 FPS and 43.48 FPS separately.
In addition, it can be observed from Table 1 that BEVFormer and PETR do not converge well in the unified detection setting while behaving promisingly trained with outdoor datasets. This phenomenon implies the difficulty of unifying indoor and outdoor 3D object detection. Through analysis, we find that BEVFormer obtains poor results when using data of all domains because its convergence is quite unstable, and the loss curve often boosts to a high value during training. PETR does not behave well since it implicitly learns the correspondence relation between 2D pixels and 3D voxels. When the camera parameters keep similar across all samples in a dataset like nuScenes [6], PETR converges smoothly. Nevertheless, when trained in a dataset with dramatically changing camera parameters like Omni3D, PETR becomes much more difficult to train.
@@ 4.2 Ablation Studies
Table 3: Ablation study on proposed strategies, which verify the effects of proposal head (PH), uneven BEV grid (UBG), sparse BEV feature projection(SBFP), and unified domain alignment (UDA). The last column presents the improvement of each row compared with the top row. $\rm AP_{3D}^{in}$ and $\rm AP_{3D}^{out}$ reflect the indoor and outdoor detection performance, respectively. Notably, although SBFP does not boost the detection precision, it reduces the computational cost of the BEV feature projection by 82.6%.
Key component designs. We ablate the effectiveness of the proposed strategies in UniMODE, including proposal head, uneven BEV grid, sparse BEV feature projection, and unified domain alignment. The experimental results are presented in Table 3. Notably, as mentioned before, the experiments in this part are conducted in the low computing resource setting due to the limited computing resources.
@@
Table 4: Ablation study on grid size and depth bin split strategy in uneven BEV grid.
@@
Uneven BEV grid. We study the effect of BEV feature grid size and depth bin split strategy in uneven BEV grid design, and the results are presented in Table 4. When the depth bin split is uneven, we split the depth bin range following Eq. 1.
Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results in Table 4, we can find that uneven depth bin deteriorates detection performance. We speculate this is because this strategy projects more points in closer BEV grids while fewer points in farther grids, which further increases the imbalanced distribution of projection features. Additionally, comparing the $1_{\rm st}$, $3_{\rm rd}$, and $4_{\rm th}$ rows of results in Table 4, it is observed that smaller BEV grids lead to better performance. We set the BEV grid size to 1m rather than 0.5m in all the other experiments due to limited computing resources and the vast training data volume of Omni3D, i.e., if we decrease the size of BEV grids, the performance of UniMODE can be further boosted compared with the current performance.
@@
It can be observed from Table 5 that when $\tau$ is 0, which means no feature is discarded, the best performance across all rows is arrived. When we set $\tau$ to $1e^{-3}$, about 82.6% of the feature is discarded while the performance of the detector remains very similar to the one with $\tau=0$. This phenomenon suggests that the discarded feature is unimportant for final detection accuracy. Then, when we increase $\tau$ to $1e^{-2}$ and $1e^{-1}$, we can find that the corresponding performances drop dramatically. This observation indicates that when we discard superfluous features, the detection precision and even training stability are influenced significantly. Combining all the observations, we set $\tau$ to $1e^{-3}$ and drop $82.6\%$ of unimportant feature in UniMODE, which reduces the computational cost by $82.6\%$ while maintaining similar performance as the one without dropping features.

### v2

@@ Abstract
Realizing unified monocular 3D object detection, including both indoor and outdoor scenes, holds great importance in applications like robot navigation. However, involving various scenarios of data to train models poses challenges due to their significantly different characteristics, e.g., diverse geometry properties and heterogeneous domain distributions. To address these challenges, we build a detector based on the bird’s-eye-view (BEV) detection paradigm, where the explicit feature projection is beneficial to addressing the geometry learning ambiguity when employing multiple scenarios of data to train detectors. Then, we split the classical BEV detection architecture into two stages and propose an uneven BEV grid design to handle the convergence instability caused by the aforementioned challenges. Moreover, we develop a sparse BEV feature projection strategy to reduce computational cost and a unified domain alignment method to handle heterogeneous domains. Combining these techniques, a unified detector UniMODE is derived, which surpasses the previous state-of-the-art on the challenging Omni3D dataset (a large-scale dataset including both indoor and outdoor scenes) by 4.9% $\rm AP_{3D}$, revealing the first successful generalization of a BEV detector to unified 3D object detection.
@@ 1 Introduction
The most critical challenge in unified 3D object detection lies in addressing the distinct characteristics of different scenarios. For example, indoor objects are smaller and closer in proximity, while outdoor detection needs to cover a vast perception range. Recently, Cube RCNN [5] has served as a predecessor in studying this problem. It directly produces 3D box predictions in the camera view and adopts a depth decoupling strategy to tackle the domain gap among scenes. However, we observe that it suffers serious convergence difficulty and is prone to collapsing during training.
@@
First of all, as shown in Fig. 1 (a) and (b), the geometry properties (e.g., perception ranges, target positions) between indoor and outdoor scenes are diverse. Specifically, indoor objects are typically a few meters away from the camera, while outdoor targets can be more than 100m away. Since a unified BEV detector needs to recognize objects in all scenarios, the BEV feature has to cover the maximum possible perception range. Meanwhile, as indoor objects are often small, the BEV grid resolution for indoor detection should be precise. All these characteristics lead to unstable convergence and significant computational burden. To address these challenges, we develop a two-stage detection architecture. In this architecture, the first stage produces initial target position estimation, and the second stage locates targets using this estimation as priori, which helps stabilize the convergence process. Moreover, we introduce an uneven BEV grid split strategy that expands the BEV space range while maintaining a manageable BEV grid size. Furthermore, a sparse BEV feature projection strategy is developed to reduce the projection computational cost by 82.6%.
@@
Combining all these innovative techniques, a Unified Monocular Object DEtector named UniMODE is developed, and it achieves state-of-the-art (SOTA) performance on the Omni3D benchmark. In the unified detection setting, UniMODE surpasses the SOTA detector, Cube RCNN, by an impressive 4.9% in terms of $\rm AP_{3D}$ (average precision based on 3D intersection over union). Furthermore, when evaluated in indoor and outdoor detection settings individually, UniMODE outperforms Cube RCNN by 11.9% and 9.1%, respectively. This work represents a pioneering effort to explore the generalization of BEV detection architectures to unified 3D object detection. It showcases the immense potential of BEV detection across a broad spectrum of scenarios and underscores the versatility of this technology.
@@ 2 Related Work
BEV detectors, on the other hand, transform image features from the 2D camera plane to the 3D physical space before generating results in 3D [12]. This approach benefits downstream tasks, as planning is also performed in the 3D space [18]. However, the challenge with BEV detectors is that the feature transformation process relies on accurate depth estimation, which can be challenging to achieve with only camera images [23]. As a result, convergence becomes unstable when dealing with diverse data scenarios [5].
@@ 3.1 Overall Framework
The feature head and depth head are responsible for projecting the image feature into the BEV plane and obtaining the BEV feature. During this projection, we develop a technique to remove unnecessary projection points, which reduces the computing burden by about 82.6% (refer to Section 3.4). Besides, we propose the uneven BEV feature (refer to Section 3.3), which means the BEV grids closer to the camera enjoy more precise resolution, and the grids farther to the camera cover broader perception areas. This design well balances the grid size contradiction between indoor detection and outdoor detection without extra memory burden.
@@ 3.2 Two-Stage Detection Architecture
The geometry property difference is identified as an essential reason causing the unstable convergence of BEV detectors [15]. For example, the difference in target position distribution makes it challenging for Transformer-based detectors to learn how to update the query reference points gradually toward the concerned objects. In fact, through visualization, we find the reference point updating in the 6 Transformer decoders is disordered. As a result, if we adopt the classical deformable DETR architecture [41] to build a 3D object detector, the training is easy to collapse due to the inaccurate positions of learned reference points, resulting in sudden gradient vanishing or exploding.
@@
To overcome this challenge, we construct UniMODE in a two-stage detection fashion. In the first stage, we design a CenterNet [39] style head (the proposal head in Fig. 2) to produce detection proposals. Specifically, its predicted attributes include the 2D center Gaussian heatmap, offset from 2D centers to 3D centers, and 3D center depths of targets. The 3D center coordinates of proposals can be derived from these predicted attributes. Then, the proposals with top $M$ confidences are selected and encoded as $M$ proposal queries by an MLP layer. To account for any potential missed targets, another $N$ randomly initialized queries are concatenated with these proposal queries to perform information interaction in the 6 decoders of the second stage (the Transformer stage). In this way, the initial query reference points of the second detection stage are adjusted adaptively. Our experiments reveal that this two-stage architecture is essential for stable convergence.
@@ 3.3 Uneven BEV Grid
A notable difference between indoor and outdoor 3D object detection lies in the geometry information (e.g., scale, proximity) of objects to the camera during data collection. Indoor environments typically feature smaller objects located closer to the camera, whereas outdoor environments involve larger objects positioned at greater distances. Furthermore, outdoor 3D object detectors must account for a wider perceptual range of the environment. Consequently, existing indoor 3D object detectors typically use smaller voxel or pillar sizes. For instance, the voxel size of CAGroup3D [31], a SOTA indoor 3D object detector, is 0.04 meters, and the maximum target depth in the SUN-RGBD dataset [29], a classic indoor dataset, is approximately 8 meters. In contrast, outdoor datasets exhibit much larger perception ranges. For example, the commonly used outdoor detection dataset KITTI [8] has a maximum depth range of 100 meters. Due to this vast perception range and limited computing resources, outdoor detectors employ larger BEV grid sizes, e.g., the BEV grid size in BEVDepth [11], a state-of-the-art outdoor 3D object detector, is 0.8 meters.
@@ 3.5 Unified Domain Alignment
Class alignment loss. In the loss view, we aim to address the heterogeneous label conflict when combining multiple data sources. Specifically, there are 6 independently labeled sub-datasets in Omni3D, and their label spaces are different. For example, as presented in Fig. 4, although the Window class is annotated in ARKitScenes, it is not labeled in Hypersim. As the label space of Omni3D is the union of all classes in all subsets, the unlabeled window in Fig. 4 (a) becomes a missing target that harms convergence stability.
@@
The two-stage detection architecture described in Section 3.2 can alleviate the aforementioned problem to some extent, because it helps the detector concentrate on foreground objects, and the unlabeled objects are overlooked to compute loss. To address this problem further, we devise a simple strategy, i.e., the class alignment loss. Specifically, denoting the label space of the $i_{\rm th}$ dataset as $\Omega_{i}$, we compute loss on the $i_{\rm th}$ dataset as:
@@ 4 Experiment
Implementation details. The perception ranges in the X-axis, Y-axis, and Z-axis of the camera coordinate system are $(-30,30)$, $(-40,40)$, $(0,80)$ meters, respectively. If without a special statement, the BEV grid resolution is $(60,80)$. The factor $\gamma$ defined in the class alignment loss is set to 0.2. $M$ and $N$ are set to 100. The adopted optimizer is AdamW, and the learning rate is set to $12e^{-4}$ for a batch size of 192. The experiments are primarily conducted on 4 A100 GPUs.
Dataset. The experiments in this section are performed on Omni3D [5], the sole large-scale 3D object detection benchmark encompassing both indoor and outdoor scenes. Omni3D is built upon six well-known datasets including KITTI [8], SUN-RGBD [29], ARKitScenes [3], Objectron [1], nuScenes [6], and Hypersim [27]. Among these datasets, KITTI and nuScenes focus on urban driving scenes, which are real-world outdoor scenarios. SUN-RGBD, ARKitScenes, and Objectron primarily pertain to real-world indoor environments. Compared with outdoor datasets, the required perception ranges of indoor datasets are smaller, and the object categories are more diverse. Hypersim, distinct from the aforementioned five datasets, is a virtually synthesized dataset. Thus, Hypersim allows for the annotation of object classes that are challenging to label in real scenes, such as transparent objects (e.g., windows) and very thin objects (e.g., carpets). The Omni3D dataset comprises a total of 98 object categories and 3 million 3D box annotations, spanning 234,000 images. The evaluation metric is $\rm AP_{3D}$, which reflects the 3D Intersection over Union (IoU) between 3D box prediction and label.
@@ 4.1 Performance Comparison
According to the results, we can observe that UniMODE achieves the best results in all metrics. It surpasses the SOTA Cube RCNN by 4.9% given the primary metric $\rm AP_{3D}$. Besides DLA34, we also try another backbone ConvNext-Base. This is because previous papers suggest that DLA34 is commonly used in camera-view detectors like Cube RCNN but unsuitable for BEV detectors [13]. Since UniMODE is a BEV detector, testing its performance with only DLA34 is unfair. Thus, we also test UniMODE with ConvNext-Base, and the result suggests that the performance is boosted significantly. Additionally, the speed of UniMODE is also promising. Test on 1 A100 GPU, the inference speeds of UniMODE under the high and low computing resource settings are 21.41 FPS and 43.48 FPS separately. Moreover, it can be observed from Table 1 that BEVFormer and PETR do not converge well in unified detection while behaving promisingly trained with outdoor datasets. This phenomenon implies the difficulty of unifying indoor and outdoor 3D object detection. Through analysis, we find that BEVFormer obtains poor results when using data of all domains because its convergence is quite unstable, and the loss curve often boosts to a high value during training. PETR does not behave well since it implicitly learns the correspondence relation between 2D pixels and 3D voxels. When the camera parameters keep similar across all samples in a dataset like nuScenes [6], PETR converges smoothly. Nevertheless, when trained in a dataset with dramatically changing camera parameters like Omni3D, PETR becomes much more difficult to train.
@@ 4.2 Ablation Studies
Table 3: Ablation studies on proposed strategies, which verify the effects of proposal head (PH), uneven BEV grid (UBG), sparse BEV feature projection(SBFP), and unified domain alignment (UDA). The last column presents the improvement of each row compared with the top row. $\rm AP_{3D}^{in}$ and $\rm AP_{3D}^{out}$ reflect the indoor and outdoor detection performance, respectively. Notably, although SBFP does not boost detection precision, it reduces the computational cost of the BEV feature projection by 82.6%.
Key component designs. We ablate the effectiveness of the proposed strategies, including proposal head, uneven BEV grid, sparse BEV feature projection, and unified domain alignment, and present the results in Table 3.
@@
Table 4: Ablation studies on grid size and depth bin split strategy in uneven BEV grid.
@@
Uneven BEV grid. We study the effect of BEV feature grid size and depth bin split strategy in uneven BEV grid design, and the results are presented in Table 4. When the depth bin split is uneven, we split the depth bin range following Eq. 1. Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results in Table 4, we can find that uneven depth bin deteriorates detection performance. We speculate this is because this strategy projects more points in closer BEV grids while fewer points in farther grids, which further increases the imbalanced distribution of projection features. Additionally, comparing the $1_{\rm st}$, $3_{\rm rd}$, and $4_{\rm th}$ rows of results in Table 4, it is observed that smaller BEV grids lead to better performance.
@@
It can be observed from Table 5 that when $\tau$ is 0, which means no feature is discarded, the best performance across all rows is arrived. When we set $\tau$ to $1e^{-3}$, about 82.6% of the feature is discarded while the performance of the detector remains very similar to the one with $\tau=0$. This phenomenon suggests that the discarded feature is unimportant for final detection accuracy. Then, when we increase $\tau$ to $1e^{-2}$ and $1e^{-1}$, we can find that the corresponding performances drop dramatically. This observation indicates that when we discard superfluous features, the detection precision and even training stability are influenced significantly. Combining all the observations, we set $\tau$ to $1e^{-3}$ and drop $82.6\%$ of unimportant features in UniMODE, which reduces the computational cost by $82.6\%$ while maintaining similar performance as the one without dropping features.